OpenStudy (anonymous):

Given: F(x) = x^2, find [F(x+h)-F(x)]/h and simplify. When I did it this is how I attempted to solve it.... [F(x) + F(h) -F(x)]/h the F(x)'s cancel.. F(h)/h The closest answer to that was just plain h, so I chose that. But it is incorrect. Can someone walk me through the process? Thanks :)

Parth (parthkohli):

Hmm... the F(x) just can't cancel... but don't worry! You'd get it right in a minute!

hartnn (hartnn):

u cannot write f(x+h) as f(x)+f(h) ...!!

Parth (parthkohli):

\[F(x)=x^2 \implies F(\text{some number}) = \text{that number}^2 \]

Parth (parthkohli):

\[F(x) = x^2\implies F(x + h) = (x + h)^2 \]Do you agree?

OpenStudy (anonymous):

Hmm. Wouldn't it be F(x^2+h) though? Because only F(x) is x^2. & It doesn't say that h is changed to anything.

Parth (parthkohli):

That's what a function means, @theequestrian.

OpenStudy (anonymous):

Wait, so it is F(x+h)^2 ? Sorry, I'm just confused.. But thanks for helping !

hartnn (hartnn):

the procedure: to find f(x+h) , u just replace x in f(x) by (x+h). remember this. ok?

OpenStudy (anonymous):

Okay, I'm with you

hartnn (hartnn):

so now replace x with (x+h) in \(x^2\) what u get??

OpenStudy (anonymous):

(x+h)^2

hartnn (hartnn):

thats your f(x+h) ok? so f(x+h)-f(x)=(x+h)^2-x^2 ok?

OpenStudy (anonymous):

Okay, then (x+h) simplifies to x^2+2xh+h^2 right?

hartnn (hartnn):

yup, so whats x^2+2xh+h^2-x^2 ?

OpenStudy (anonymous):

2xh+h^2

hartnn (hartnn):

great :) now the question was [F(x+h)-F(x)]/h u got numerator as 2xh+h^2 so whats (2xh+h^2)/h = ?

OpenStudy (anonymous):

2x+h ?

hartnn (hartnn):

thats your final answer.. good work :)

OpenStudy (anonymous):

One question. I thought you couldn't reduce with addition or subtraction in the fraction?

hartnn (hartnn):

i did not understand your question...can u give example?

OpenStudy (anonymous):

Like in (2xh+h^2)/h there is addition in the numerator. I thought you could not cancel out anything from the bottom if there was addition or subtraction (addition in this case) in the numerator ? Does that make more sense?

hartnn (hartnn):

\[\frac{2xh+h^2}{h}=\frac{h(2x+h)}{h}=2x+h\] ok?

OpenStudy (anonymous):

Ohhh. I see now! Thanks for clarifying :)

hartnn (hartnn):

welcome :)